OpenStudy (anonymous):

If the discriminant of a quadratic equation is zero, determine the number of real solutions? a. no real solutions b. one real solution c. two real solution

OpenStudy (boldjon):

A

OpenStudy (anonymous):

thanks

OpenStudy (boldjon):

np

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x=\frac{-b\pm\sqrt{\color{red}{~0}~}}{2a}=\frac{-b}{2a} }\) So that means that how many solutions are there when the discriminant is zero?